OpenStudy (anonymous):

What is the product of the product of 3 times the square root of 6 x and 2 times the square root of 15 x? Simplify if possible.

OpenStudy (anonymous):

3(6x^2)=2(15x^2)

OpenStudy (lgbasallote):

\[\large (3 \sqrt{6x}) \times (2 \sqrt{15x}) = (3\times 2) \times (\sqrt{6x \times 15x})\]

OpenStudy (anonymous):

18x?

OpenStudy (lgbasallote):

therefore you get \[\large 6 \sqrt{90x^2}\] simplifying that you get \[\large 6(3)(x) \sqrt{10} = 18x\sqrt{10}\]

OpenStudy (anonymous):

Thanks Igbasallote!